OpenStudy (anonymous):

The value of a graphing calculator is $225. After 2 years, the value of this calculator is $160. Find the value of the calculator after 5 years

OpenStudy (anonymous):

assuming it goes down in value exponentially it would be \[225\times (\frac{160}{225})^\frac{5}{2}\]